OpenStudy (anonymous):

solve each literal equation for the given variable Solve M = 2x + 3y for y

OpenStudy (swissgirl):

ummm I am guess we wld follow what angela did

OpenStudy (anonymous):

thats what im trying to go off of

OpenStudy (swissgirl):

\( M=2x+3y \) \(M-2x=3y\) \( \frac{M-2x} {3} =y \)

OpenStudy (swissgirl):

can you read what i did its a lil small

OpenStudy (anonymous):

yeah i see it do you mind like explaining to me how you did it like steps and such so i can learn to do it on my own?

OpenStudy (swissgirl):

okk well you see we started off with the original equation

OpenStudy (swissgirl):

now we need to try to isolate y so first I brought over the 2x. So i subtracted 2x from both sides M-2x=2x+3y-2x M-2x=3y

OpenStudy (swissgirl):

and then i divided 3 from both sides \( \huge\frac{M-2x} {3} = \frac{3y} {3} \) \( \huge\frac{M-2x} {3} =y \)

OpenStudy (swissgirl):

and so Voila we have isolated our y

OpenStudy (swissgirl):

idk if u followed but i gotta go work I am nutssss

OpenStudy (anonymous):

im kinda getting it better thank you :)